OpenStudy (anonymous):

what is the slope of the tangent line to the curve (x^3 + y)^2 -2a^2xy = b^2 at the point P(0,1) when a=(2)^(1/2) and b = 1

OpenStudy (anonymous):

@ganeshie8 @Robert136 @nincompoop

OpenStudy (anonymous):

I'm getting DNE but its wrong

OpenStudy (jhannybean):

For this problem can you input the values for a and b, then find the derivative? Or does the inputting of values happen after differentiating..

OpenStudy (jhannybean):

It looks really ugly if you differentiate it along with a and b :\

OpenStudy (jhannybean):

I would use implicit differentiation to find the derivative.

OpenStudy (anonymous):

i really have no idea , iv imputed the values for a and be and then solved for y and derived that , but that seems to be wrong

OpenStudy (jhannybean):

Yeah it is hahahahaha

OpenStudy (jhannybean):

Because that's basically stating the tangent line is found at a point before even being found.

OpenStudy (anonymous):

you only need to find f'(x,y)

OpenStudy (anonymous):

but before u start it says when a=(2)^(1/2) and b = 1 , a and b seems like small values and you can sub them why it mention them in ur question ? seems like there is a formula in ur book or something

OpenStudy (irishboy123):

\[\vec n = <2(x^3+y)3x^2 - 2a^2y, 2(x^3+y)-2a^2x>\] \[= <6x^2(x^3+y) - 2a^2y, 2(x^3+y)-2a^2x>\] x = 0 \[= <-2a^2y,2>\] \[y = 1, a = \sqrt{2}\] \[\vec n = <-4,2,>\]

OpenStudy (freckles):

don't solve for y just find the derivative of both sides w.r.t. x a and b are constants you can input those values in later or not once you have differentiated both sides you can input the values for a and b if you haven't already and then also input 0 for x and 1 for y then solve for y' to find the slope.

OpenStudy (anonymous):

see thats what i'm talking about :D like @irishboy123

OpenStudy (irishboy123):